# MIGHTY: Hermite Spline-based Efficient Trajectory Planning

### v3

@@ IV-G Dynamic Environments
We further evaluate MIGHTY in two dynamic scenarios: (1) only dynamic obstacles, and (2) dynamic & static obstacles. For both, in addition to the cost terms described in Section IV-F, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$. We set the weight $w_{\text{dyn}}=${10}^{1}$\$ and use a soft barrier radius $C_{\text{dyn}}=$3.0\text{\,}\mathrm{m}$\$. Note that $C_{\text{dyn}}$ is where the penalty activates, not a hard safety distance. The robot collision radius is $0.1\text{\,}\mathrm{m}$. The start and goal are $(0,0,3)\,$\mathrm{m}$\$ and $(100,0,3)\,$\mathrm{m}$\$, and, in this more complex environment, the dynamic constraints are set to $v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$.
@@
Dynamic & static environment: We add 50 dynamic obstacles (same motion model) to the static forest as shown in Fig. 9 and run 10 trials. MIGHTY avoids all obstacles in all trials. The minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$.
@@ V Hardware Experiments
We also evaluated MIGHTY in three hardware experiment: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission. The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [2]. Planning runs onboard on an Intel NUC 13, and low-level control uses PX4 [11] on a Pixhawk flight controller [12]. All perception, planning, and control run in real time onboard.
@@
High-Speed Flight Experiment: We further evaluate performance at higher speeds. As shown in Section IV, MIGHTY tends to produce higher-jerk trajectories; when constraints are satisfied, these trajectories remain safe while enabling faster motion. We test $v_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$6.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Figure 11 shows the resulting histories. All flights completed without collision, demonstrating that the MIGHTY planner and DLIO state estimation operate effectively at high speed. At $v_{\max}=$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, the vehicle reached $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ peak speed.
@@
Dynamic Obstacle Flight Experiment: To assess robustness to changes in the environment, we introduce and move obstacles during the mission (Fig. 12). We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Section IV-G for details. The agent successfully reaches the goal without collision over $490\text{\,}\mathrm{s}$.

### v4

@@ IV-G Dynamic Environments
We further evaluate MIGHTY in two dynamic scenarios: (1) only dynamic obstacles, and (2) dynamic & static obstacles. For both, in addition to the cost terms described in Sec. IV-F, we add a dynamic-obstacle avoidance cost $J_{\text{dyn}}=\int_{0}^{\sum_{s}T_{s}}\big(C_{\text{dyn}}^{2}-\|\mathbf{x}(t)-\mathbf{k}(t)\|^{2}\big)_{+}^{3}\,dt.$ where $\mathbf{k}(t)$ is the position of a dynamic obstacle at time $t$, and $(\cdot)_{+}=\max(0,\cdot)$. We set the weight $w_{\text{dyn}}=${10}^{1}$\$ and use a soft barrier radius $C_{\text{dyn}}=$3.0\text{\,}\mathrm{m}$\$. Note that $C_{\text{dyn}}$ is where the penalty activates, not a hard safety distance. The robot collision radius is $0.1\text{\,}\mathrm{m}$. The start and goal are $(0,0,3)\,$\mathrm{m}$\$ and $(100,0,3)\,$\mathrm{m}$\$, and, in this more complex environment, the dynamic constraints are set to $v_{\max}=$2.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, $a_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{2}$\$, and $j_{\max}=$30.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}^{3}$\$.
@@
Dynamic & static environment: We add 50 dynamic obstacles (same motion model) to the static forest as shown in Fig. 9 and run 10 trials. MIGHTY avoids all obstacles in all trials. The minimum is $1.0\text{\,}\mathrm{m}$, above the $0.1\text{\,}\mathrm{m}$ collision radius.
@@ V Hardware Experiments
We also evaluated MIGHTY in three hardware experiments: (1) long-duration flights at $v_{\max}\in\{1,2,3,4\}\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, (2) high-speed flights up to $v_{\max}=7\,$\mathrm{m}\mathrm{/}\mathrm{s}$\$, and (3) flights with obstacles introduced during the mission. The perception uses a Livox Mid-360 LiDAR, and localization is provided by DLIO [25]. Planning runs onboard on an Intel NUC 13, and low-level control uses PX4 [26] on a Pixhawk flight controller [27]. All perception, planning, and control run in real time onboard.
@@
High-Speed Flight Experiment: We further evaluate performance at higher speeds. As shown in Sec. IV, MIGHTY tends to produce higher-jerk trajectories; when constraints are satisfied, these trajectories remain safe while enabling faster motion. We test $v_{\max}=$5.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$6.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$,$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$. Fig. 11 shows the resulting histories. All flights completed without collision, demonstrating that the MIGHTY planner and DLIO state estimation operate effectively at high speed. At $v_{\max}=$7.0\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$\$, the vehicle reached $6.7\text{\,}\mathrm{m}\mathrm{/}\mathrm{s}$ peak speed.
@@
Dynamic Obstacle Flight Experiment: To assess robustness to changes in the environment, we introduce and move obstacles during the mission (Fig. 12). We track a person carrying obstacles and incorporate the estimates into the trajectory optimization; see Sec. IV-G for details. The agent successfully reaches the goal without collision over $490\text{\,}\mathrm{s}$.
@@ VI Conclusions
VII Acknowledgements
The authors would like to thank Lili Sun for her help with the hardware experiments and Mason B. Peterson for his support with robotdatapy library. The authors also thank Juan Rached for his help with hardware setup. This work was supported in part by DSTA and ARL grant W911NF-21-2-0150.